OpenStudy (anonymous):

log2^3+log2^x=log2^20

OpenStudy (raden):

simplify for left side : log2^3+log2^x=log2^20 log(2^3 * log2^x)=log2^20 log2^(3+x)=log2^20 now solve for x 3+x=20

OpenStudy (anonymous):

Thank You!

OpenStudy (raden):

welcome